### v2

# Multi-objective Representation for Numbers in Clinical Narratives Using CamemBERT-bio

Abstract
The processing of numerical values is a rapidly developing area in the field of Large Language Models (LLMs). Despite numerous advancements achieved by previous research, significant challenges persist, particularly within the healthcare domain. This paper investigates the limitations of LLMs in understanding numerical values. Objective: The aim of this research is to categorize numerical values extracted from medical documents into eight specific physiological categories using CamemBERT-bio. Methods: Building on prior findings that reveal potential limitations of transformer-based models, we examine two strategies: fine-tuning CamemBERT-bio on a small medical dataset with the integration of Label Embedding for Self-Attention (LESA), and combining LESA with additional enhancement techniques such as Xval. Given that CamemBERT-bio is already pre-trained on a large medical dataset, the first approach aims to update its encoder with newly added label embeddings technique, while the second approach seeks to develop multiple representations of numbers (contextual and magnitude-based) to achieve more robust number embeddings. Results: As anticipated, fine-tuning the standard CamemBERT-bio on our small medical dataset did not improve F1 scores. However, significant improvements were observed with CamemBERT-bio + LESA, resulting in an over 15% increase. Similar enhancements were noted when combining LESA with Xval, outperforming conventional and previous methods. Conclusions and Novelty: This study introduces two innovative techniques for handling numerical data, which are also applicable to other modalities. We illustrate how these techniques can improve the performance of Transformer-based models, achieving more reliable classification results even with small datasets.
Index Terms:
Clinical Natural Language Processing, Numerical values classification, Language models training.
Clinical and Translational Impact Statement— The use of CamemBERT-bio to numerical value classification holds significant promise, particularly given the limited availability of clinical data. By developing multiple representation-based embeddings for numbers, we enhance our models robustness and generalizability. This approach enables the creation of diagnostic algorithms from unstructured medical notes, as our model provides the necessary structure. The methodology can be extended to various other NLP challenges and modalities, leveraging the core idea of constructing embeddings based on multiple data representations. Additionally, this work addresses the common healthcare issue of handling small and imbalanced datasets.
I Introduction
A major challenge in processing medical texts is understanding numerical data. Numbers are ubiquitous in medical documents, appearing in various forms such as physiological measurements, DNA coding sequences, event occurrences (e.g., number of births, cardiac failures), durations (e.g., pregnancy length, medication duration), and medical protocol codes (e.g., G1P2 standing for Gravida 1 Para 2). In the context of Clinical Decision Support Systems (CDSS), it is crucial to recognize and interpret the meaning of these numbers within their specific context to determine if they indicate a potential problem. For example, in the text ”Patient agé de 12 ans, rythme cardiaque de 130 bpm,” it is essential to identify ”12” as the age of the patient and ”130” as the heart rate, and then understand that a heart rate of ”130” is problematic for a 12-year-old patient.
Most machine learning efforts in the medical domain must contend with small and imbalanced datasets due to the high cost of data annotation in healthcare. Additionally, medical notes are often unstructured, filled with medical jargon, and sometimes contain incomplete sentences. This necessitates handling shorthand notations, typographical errors, and various forms of textual noise commonly found in medical documentation, making systematic approaches like case-by-case algorithms ineffective. In addition, numerical values in medical datasets typically span a wide range. For example, the dataset provided by Centre Hospitalier Universitaire Sainte-Justine (CHUSJ) for our study includes numerical values ranging from $10^{-4}$ to $10^{5}$. Therefore, careful gradient management is essential, as a single large outlier can negatively impact the overall performance of the model.
This paper contributes to a broader project aimed at leverage Pretrained Large Language Models (PLMs) for diagnosing cardiac failure. According to the National Heart, Lung, and Blood Institute (2023), cardiac failure occurs when the heart is unable to pump sufficient blood to meet the body’s needs, either due to inadequate filling or diminished pumping capacity. Cardiac failure presents through a range of physiological indicators, including blood pressure, ventricular gradients, and ejection fractions. Our research focuses on classifying numerical values extracted from medical notes into one of eight specific physiological categories using CamemBERT-bio [1].
I-A Goal statement
Our objective is to train CamemBERT-bio on a limited and uneven dataset with the aim of accomplishing the following outcome:
Input medical note:
“Norah, BB né à 37+6 par AVS, APGAR 8-8-8. Dx Anténatale: d-TGV avec septum intact, sat 50-65% à la naissance.”
Expected output:
Value Attributes Unit
$37+6$ Divers (âge du patient) semaines
$8-8-$ 8 - 8 score APGAR
[…]
I-B Contribution
In this work, we outline the subsequent contributions:
  • We demonstrate that CamemBERT-bio + LESA as introduced in [2] can be further enhanced through language modeling prefinetuning. This prefinetuning does not improve CamemBERT-bio alone, indicating that integrating the Label Embedding for Self-Attention (LESA) technique necessitates pretraining to adjust the model’s parameters. Prefinetuning on a small medical dataset significantly improves the F1 score for CamemBERT-bio + LESA.
  • We develop a new model by extending finetuned CamemBERT-bio + LESA with the Xval architecture [3]. The main idea of Xval is to incorporate the magnitude of numbers into their embeddings before the attention layers. This new model provides a multi-objective representation of numbers, enhancing generalizability and robustness.
  • We assess and juxtapose our tailored model against conventional counterparts. Findings demonstrate that our model surpasses standard approaches and adeptly manages datasets characterized by limited size and imbalance.
II Related works
Processing numerical values with LLMs is a highly active research area. While [4] claimed that most word embedding models could capture numeracy, several limitations persist. Various approaches have been developed to handle numerical values. [5] tackled the issue of number classification by employing a rules-based method utilizing regular expressions. Nevertheless, the success of this approach is largely contingent on the clarity of the text. It depends significantly on consistent abbreviations and the minimal presence of irrelevant textual noise, which can disrupt the regular expression generator’s accuracy. [6] proposed a question-answer pre-training task aimed at improving number understanding. Other studies have explored the potential of different number tokenization methods, such as digit-by-digit, scientific notation, and customized notations [7]. Some previous studies have tackled the challenge with a number-agnostic approach, replacing numbers with placeholder keywords and scaling their embeddings to enforce context-centric encoding [3, 8, 2]. In their work, [2] introduced Label Embedding for Self-Attention (LESA), which incorporates class descriptions directly into the initial feature extraction phase to help the feature extractor focus on relevant features for classification. LESA is similar to EmbedNum [9], which classifies numbers by comparing their embeddings to those of potential categories. However, LESA differs by using similarity exclusively to construct the number’s embedding rather than for the classification process itself.
Another significant challenge is handling small and imbalanced datasets. While transformer-based architectures generally perform well with large datasets, research by [10, 11, 12] indicates that they may not outperform traditional methods, such as LSTM, when dealing with smaller datasets of clinical notes. A common technique to enhance LLM performance on small datasets is Knowledge Distillation (KD) [13]. Knowledge Distillation (KD) is based on the principle that a model with fewer parameters is less prone to overfitting. This technique entails training a smaller neural network to replicate the performance of a larger one, using a teacher-student framework. This methodology is the key idea of like DistilBERT [14] and its French counterpart, DistillCamemBERT [15]. Another approach to handling limited data is to create multiple representations for words, leveraging diverse methods to achieve more robust and generalizable modeling. Typically, this involves constructing an objective loss as the sum of various losses, each specialized in different aspects of the studied modality. However, this approach raises the critical issue of appropriately weighting each loss to avoid bias. In [16], the authors propose an automatic method for weighting multiple losses by optimizing the log-likelihood of the model predictions. Although their approach was applied to images, we found their results relevant and potentially beneficial for our case.
III The task and the dataset
The dataset utilized in this study was supplied by the Pediatric Intensive Care Unit at CHU Sainte-Justine (CHUSJ), and includes children under the age of 18. It comprises 26,671 unannotated medical notes from 2,514 de-identified patients. Of these notes, only 1,072 samples (approximately 30,000 tokens) are annotated. All the details regarding the dataset are provided in [2].
The number distribution is displayed in Figure  1. All the numbers beyond $500$ are regrouped because they are scattered from $500$ to $35000$.
Figure 1: (a) Overlook to the whole numbers distribution (b) A focus on the small numbers distribution. Non integers numbers are effectively represented in the dataset.
IV A few limitations of end-to-end learning for LLM in medical datasets
End-to-end learning involves training a model to perform a task from raw input to final output, without intermediate steps or feature engineering. While this is the standard approach for training most LLMs, our work faces the significant limitation of working with small and imbalanced datasets, unlike many state-of-the-art approaches mentioned earlier. Additionally, there are several other major limitations of end-to-end training for LLMs in the medical domain.
IV-A The limitations of Tokenization
A fundamental aspect of recent transformer-based models revolves around the introduction of highly efficient tokenization techniques, notably exemplified by WordPiece [17] and SentencePiece [18]. These tokenizer variants share a common trait: the capability to deconstruct intricate or infrequent words into subsequences comprising their constituent characters, an important ability in managing Out-of-vocabulary (OOV) words. Let’s consider the word “working”, which undergoes tokenization into “work” + “ing”. This resultant sequence forms a good starting point to model the word, as “work” and “ing” independently hold semantic significance and individual meaning. However, as mentionned by [8], we claim that this approach exhibits limitations when dealing with numerical values. For instance, consider the instance of “123.7g/L”, which would be disassembled into “12” + “3” + . + “7” + “g/L”. This segmentation raises several issues. Predominantly, in many classification tasks, only the first token of a segmented word is subject to classification. Only “12” would be classified in our example. However, “12” is markedly ambiguous, vastly deviating from the initial term’s inherent meaning. Furthermore, this method of breaking down numerical tokens might fail to capture their magnitudes. The difference in magnitude between “12” and “123.7” is substancial. While the self-attention mechanism and index position embedding hold the potential to discern this magnitude discrepancy, we claim that these mechanisms alone are insufficient.
Figure 2: Output of a completion task by CamemBERT-bio. The input sentence is “Patient en détresse respiratoire, gradient VG-VD ad $\langle$ mask $\rangle$ mmgh.”. The figure contains the possible words to fill the mask with their corresponding probabilities.
[…]
IV-B The structure issue
Some earlier research focused on studying transformer-based models to understand how they function and evaluate their ability to recognize different syntactic relations. [19] revealed that, surprisingly, without explicit supervision on syntax and co-reference, certain heads in the BERT model specialize in grammatical tasks such as finding direct objects of verbs, determiners of nouns, objects of prepositions, and objects of possessive pronouns with more than 75% accuracy. This reveals that the structure of sentences matters for these models. However, our medical notes landscape stands apart in this regard. A significant proportion of sentences within these notes lack well-defined structures, with instances wherein clear subjects and verbs are even absent, as aptly exemplified by the following illustrations:
  • CARDIOMYOPATHIE 1) S’est présenté le 28/04 à l’hôpital - Admis ad 07/05 aux soins […] Depuis, céphalée on/off, fatigue, essoufflement.[…] Vomissement le soirx 5 jours.eeeee Pincement thoracique à 2-3 reprises depuis 1mois. l’a réveillé 1 fois.
  • ATCDp: #PNA droite à 1 mois de vie et en 2004 eeeee Atrophie/asymétrie rénale 2aire Vu en néphro: eeeee Écho (11/2014): Rein Droit: 8.8 cm, contour supérieur légèrement irrégulier + vague zone kystique, de 5 mm, […].
  • “[…] FR 21 FC 100-110 FR 50 […].”
As noted in [2], the last excerpt underscores the lack of contextual information in sentence structures. This excerpt reads a sequence of physiological indicators. The initial reference to FR most likely indicates the respiratory rate (fréquence respiratoire) since a value of 21 is usually too low for the shortening fraction (fraction de raccourcissement). Conversely, the second occurrence of FR refers to a value above 50, suggesting it pertains to the shortening fraction. In this case, it was the values themselves, rather than the context, that provided the necessary distinctions, highlighting a flaw in the sentence structure. This issue echoes the challenges of tokenization, emphasizing the need for a thorough understanding of the numerical values. Additionally, It’s worth mentioning that CamemBERT-bio underwent no pre-training using unclear or poorly structured medical documents.
V Methodology
The methodology adopted in this study involves developing a multi-objective representation of numbers to achieve a more generalizable and robust model [20]. Our objectives are Masked Language Modelling (MLM) loss and Number Prediction Loss. These two objectives result in the derivation of two models:
  • CamemBERT-bio + LESA prefinetuned via MLM on our small medical dataset
  • CamemBERT-bio + LESA + Xval trained with both MLM loss and Number Prediction Loss.
We anticipate that this multi-objective training approach will produce number representations that incorporate both contextual and magnitude information.
V-A Masked Language Modelling (MLM) objective
We summarize the model description from [2]. Let’s consider a text $\boldsymbol{t}=[token_{1},token_{2},\cdots,token_{L}]$ of $L$ tokens, with initial embeddings $[\boldsymbol{x}_{token_{1}},\cdots,\boldsymbol{x}_{token_{L}}]$. These initial embeddings, prefixed with the special token $[CLS]$, are fed into the transformer layers in the following manner:
$\boldsymbol{X}=[\boldsymbol{x}_{CLS},\boldsymbol{x}_{token_{1}},\cdots,\boldsymbol{x}_{token_{L}}]\in\mathbb{R}^{(L+1)\times D}$ (1)
where $D$ is the dimension of the embedding space.
For every individual class, some descriptive keywords are selected, and their initial embeddings are computed and then averaged. This calculation produces an embedding matrix, represented as $\boldsymbol{X}^{l}\in\mathbb{R}^{n\times D}$, where $n$ corresponds to the number of classes. This matrix includes unique label embeddings for each class.
[…]
Then, after computing the cosine similarity matrix $\textbf{{CoSim}}=\textrm{norm}(\boldsymbol{A}^{l})^{T}\textrm{norm}(\boldsymbol{A}^{l})$, we therefore get the enhanced self-attention:
[…]
Figure 3: Excerpt from [2]. An illustration of the Label Embedding for Self-Attention (LESA) is shown. The input to this Self-Attention layer consists of the token embeddings $[X_{CLS},X_{1},\cdots,X_{L}]$ and the keyword embeddings $[X^{l}_{1},X^{l}_{2},\cdots,X^{l}_{n}]$. This layer outputs the enhanced self-attention.
In this work, we train this architecture using the Masked Language Modelling (MLM) task as a prefinetuning step. The MLM task involves randomly masking tokens and asking the model to predict the masked tokens, thereby training the model to build contextual representations of tokens. To address the limited size of our dataset, we employ the LESA mechanism to accelerate the learning process. Essentially, we inject additional information into the model via label-related keywords, enabling the model to more quickly grasp the relationships between tokens.
V-B Number Prediction objective
This method consists in four steps as detailed in [3]. Let’s consider an input text $t$ containing both numbers and words. First, all the numbers are replaced by a placeholder keyword NUM, forming a new input text $t_{masked}$. The initial numbers are saved for later. Then, we tokenize and embed $t_{masked}$ resulting in the matrix $h_{masked}$. Subsequently, all the vectors in $h_{masked}$ corresponding to the numbers previously replaced are multiplied by their corresponding value, resulting in a new matrix $h$. Finally, this embeddings matrix $h$ is fed to the encoder of our model.
By replacing all numbers by a keyword placeholder, we solve the tokenization limitation mentionned earlier. This loss objective aims at training the model to learn the magnitude of the numbers it encodes. The magnitude information is very important as it is a good discriminator between classification classes. For instance the numbers belonging to heart rate class are usually bigger than those belonging to APGAR class.
V-C Multi Objective Loss
Working with a multi objectives losses require to merge them into one single loss
$L=\sum w_{i}L_{i}$
where $w_{i}$ represents the weight of the loss $L_{i}$. In [16], the authors propose an automatic way to weight multiple losses by optimizing the model predictions log-likelihood. In plain terms, for the number prediction task, the authors assume that the likelihood is given as a Gaussian
$p\left(y|f_{1}(x,w)\right)=\mathcal{N}\left(f_{1}(x,w)|\sigma_{1}^{2}\right)$ (3)
where $x$ is the input sequence, $w$ the model parameters and $f_{1}(x,w)$ the model’s number prediction. $\mathcal{N}$ refers to the gaussian probability distribution function and $\sigma_{1}$ its standard deviation. This leads to the log likelihood
$\log p(y|f_{1}(x,w))\propto-\frac{1}{2\sigma_{1}^{2}}\left(y-f_{1}(x,w)\right)^{2}-\log\sigma_{1}$ (4)
Then one can recognize the square error $\left(y-f_{1}(x,w)\right)^{2}$ with its weight $\frac{1}{2\sigma_{1}^{2}}$ and a regularization term $\log\sigma_{1}$.
This appraoch also works for the MLM task. We write the likelihood as
$p(z|f_{2}(x,w),\sigma_{2})=\textrm{Softmax}\left(\frac{1}{\sigma_{2}^{2}}f_{2}(x,w)\right)_{z}$ (5)
where $f_{2}(x,w)$ is the predicted token distribution vector and $\sigma_{2}$ is artificially added as a temperature in a Boltzmann distribution. Then, after a few approximations, we get to write the log likelihood as
$\log p(z|f_{2}(x,w))\propto\frac{1}{\sigma_{2}^{2}}\log\textrm{Softmax}\left(f_{2}(x,w)\right)_{z}-\log\sigma_{2}$ (6)
Here, one may recognize the cross entropy loss with its weight $\frac{1}{\sigma_{2}^{2}}$ and a regularization term $\log\sigma_{2}$. When we sum up everything, for a given number $y$, we consider the following loss:
$\begin{split}\mathcal{L}(y)&=\frac{1}{2\sigma_{1}^{2}}\left(y-f_{1}(x,w)\right)^{2}-\frac{1}{\sigma_{2}^{2}}\log\textrm{Softmax}\left(f_{2}(x,w)\right)_{y}\\ &+\log\sigma_{1}+\log\sigma_{2}\\ &=\frac{1}{2\sigma_{1}^{2}}\textrm{SEL}+\frac{1}{\sigma_{2}^{2}}\textrm{MLM}_{loss}+\log\sigma_{1}+\log\sigma_{2}\end{split}$ (7)
where SEL stands for the Square Error Loss, that is $\left(y-f_{1}(x,w)\right)^{2}$, and $\textrm{MLM}_{loss}$ is the language modeling loss, that is $\log\textrm{Softmax}\left(f_{2}(x,w)\right)_{y}$.
Now we optimize this lost with respect to $w,\sigma_{1},\sigma_{2}$. In our case, when writing the first-order condition with respect to $\sigma_{1}$, we get
$\sigma_{1}=\left(y-f_{1}(x,w)\right)$ (8)
Since the exact numbers within a medical note are often not predictable from the context, we anticipate that the regression loss $(y-f_{1}(x,w))^{2}$ will have a significant lower bound that we will note $l$. Additionally, this lower bound increases with the widening range of possible values, which is particularly high in our case due to the broad spectrum of numbers. Consequently, at equilibrium, Equation  8 implies that $\sigma_{1}\approx\sqrt{l}$ and the weight of the regression loss $\frac{1}{2\sigma_{1}^{2}}\approx\frac{1}{2l}$, which is very low. Experimentally, we find that this weight is significantly lower than that of the MLM loss, placing greater emphasis on contextual representation over magnitude-based representation. As a result, we observe poor performance.
To address this issue, we equalized the weights of the two losses and substituted the regression Square Error Loss $\left(y-f_{1}(x,w)\right)^{2}$ with the Logarithmic Square Error Loss (LSEL):
$\left(\log(y+1)-\log\left(f_{1}(x,w)+1\right)\right)^{2}$ (9)
We therefore get the new multi-objective loss
$\mathcal{L}(y)=\frac{1}{2}\textrm{LSEL}+\frac{1}{2}\textrm{MLM}_{loss}$ (10)
This substitution is motivated by the following considerations. Let $\omega$ represent a parameter that requires updating. For simplicity, we’ll denote the partial derivative with respect to $\omega$ as $\partial_{\omega}$. We derive the two following gradients of SEL and LSEL with respect to $\omega$:
$\begin{split}\partial_{\omega}\textrm{SEL}\left(f_{1}(x,w),y\right)&=\partial_{\omega}\left\{\left(y-f_{1}(x,w)\right)^{2}\right\}\\ &=2\partial_{\omega}f_{1}(x,w)\left(f_{1}(x,w)-y\right)\end{split}$ (11)
$\begin{split}\partial_{\omega}\textrm{LSEL}\left(f_{1}(x,w),y\right)&=\partial_{\omega}\left\{\left(\log(y+1)-\log\left(f_{1}(x,w)+1\right)\right)^{2}\right\}\\ &=2\frac{\partial_{\omega}f_{1}(x,w)}{f_{1}(x,w)+1}\log\left(\frac{f_{1}(x,w)+1}{y+1}\right)\end{split}$ (12)
Now, the two gradients can be compared by examining their ratio.
$\frac{\partial_{\omega}\textrm{LSEL}}{\partial_{\omega}\textrm{SEL}}=\frac{1}{\left(f_{1}(x,w)+1\right)\left(f_{1}(x,w)-y\right)}\log\left(\frac{f_{1}(x,w)+1}{y+1}\right)$ (13)
Here, the term $\frac{1}{f_{1}(x,w)-y}\log\left(\frac{f_{1}(x,w)+1}{y+1}\right)$ is contingent on the relative comparison between $f_{1}(x,w)$ and $y$. However, the term $\frac{1}{f_{1}(x,w)+1}$ exhibits a dependency on the magnitude of the prediction. In simpler terms, when high numbers (outliers) are predicted, the Square Error Loss (SEL) will significantly update the weights compared to the Logarithmic Square Error Loss (LSEL), potentially undermining the model’s performance on the real data distribution. This characteristic underscores why LSEL is more effective in handling a wide range of numbers.
Figure  4 provides a concise visual representation of the multi-objective loss mechanism. In this illustration, the first masked token is decoded as the standard word ”Saturation,” assigned the default numerical value of $1$ for non-numerical tokens. The second masked token is decoded as the special token ”NUM,” corresponding to the value $65$.
Figure 4: Brief illustration of the multi objective loss.
VI Experiments
This section introduces the baseline models, outlines our experimental configuration, and presents the results we achieved.
VI-A Our models
We present two models that incorporate the diverse solutions proposed in this research.
Model 1:aaa The architecture is exactly the one presented in [2]. The training processus comprises the 2 following steps:
  1. 1.
    We pre-finetune the pre-trained CamemBERT-bio augmented with the Label Embedding for Self-Attention (LESA) technique on the unannotated medical notes with the Masked Language Modeling (MLM) task.
  2. 2.
    We then fine-tune the pre-finetuned model on the token classification task with the annotated notes.
We kept the same architecture as in [2]
Model 2:aaa We incorporated Xval architecture to Model 1, in order to have both contextual and magnitude representations of numbers.
VI-B Baselines
DistillCamemBERT:aaaDistilCamemBERT, as described in [21], employs Knowledge Distillation [13], [22] in order to reduce the parameters of CamemBERT. This technique is a training method where a larger, expert model (teacher) guides a smaller model (student) to mimic its behavior. DistillCamemBERT has achieved performance levels comparable to CamemBERT on various NLP tasks, providing a benchmark for evaluating our approach against parameter reduction techniques. For this experiment, we utilized the version available on Huggingface [23].
Camembert-bio:aaa This transformer based model is a french version of BioBERT [24] introduced by [1] and trained on a large medical corpus.
CamemBERT-bio + LESA: This model follows the training approach detailed in [2], which consists in incorporating LESA to CamemBERT-bio. The authors derived two models from this approach: one model trained on the raw dataset and another where numbers are replaced with a keyword placeholder before training. We will only keep the second one as it got the best performances. We will refer to this baseline as CamemBERT-bio + LESA
CamemBERT-bio + Xval: This model follows the training approach detailed in the work by [3], which consists in incorporating some magnitude information into the numbers embeddings. The architecture was imported from their GitHub repository.
NumBERT:aaa As intruduced in [25], it is a version of the BERT model trained by substituting each numerical instance in the training dataset with its scientific notation representation, comprising both an exponent and mantissa. We used the implementation provided in their GitHub repository.
ELMO:aaa As introduced in [26], Embeddings from Languages Models (ELMO) is a bidirectional LSTM that is trained with a coupled language model (LM) objective on a large text corpus. ELMO is very effective in practice in medical related classification tasks. We used the implementation from their GitHub repository.
VI-C Setup and Evaluation
The performance evaluation metric utilized is the $F_{1}$ score, a class-wise metric particularly suitable for imbalanced datasets. For the pre-finetuning step, we trained all our models for 100 epochs as a warm up with AdamW optimizer [27] optimizer and a learning rate of $53e-$ 5. Then we used cosine learning scheduler until early-stopping with 4 epochs of patience. For the classification downstream task, we performed 10 sets of runs with different random seeds until early stopping. For all the methods considered in this study, we used cross-entropy loss with appropriate weights for each classes in order to compensate the distribution imbalance. However, such weights were not necessary for our proposed Model 2 which performed better without them. The training, validation, and testing datasets were structured to maintain a distribution of 70%, 15%, and 15% respectively across all eight classes. Our models’ training process took 20 hours on a Tesla V100 GPU with 32 GB of memory. In contrast, training the baseline models required approximately 10 hours. The entire codebase, without the dataset and the trained model weights, can be accessed on our GitHub repository.
VII Results and Discussions
Table  III shows that all models exhibit nearly perfect $F_{1}$ scores for the Out-of-class class. This was expected given its extensive representation. To evaluate overall performance, we compute the average $F_{1}$ score across all classes without applying weighting, considering that the remaining classes hold equal importance in terms of size. This choice is made to ensure a straightforward evaluation without introducing unnecessary complexity. The overall $F_{1}$ scores of all models are presented in Table  I.
TABLE I: Overall $F_{1}$ score per model
Model Overall $\mathbf{F_{1}}$ score
ELMO 0.72
DistilCamemBERT 0.74
Camembert-bio 0.70
Camembert-bio + LESA 0.83
Camembert-bio + Xval 0.80
NumBERT 0.77
Model 1 0.83
Model 2 0.86
Using this ranking as our basis, it becomes evident that our two models (models 1 and 2) consistently surpass all the other baseline methods.
VII-1 Impact of prefinetuning via MLM
Table II presents a performance comparison of several models with and without prefinetuning. Prefinetuning has no impact on the pretrained CamemBERT-bio model, which is expected since CamemBERT-bio was trained on diverse medical datasets similar to our current dataset, except for the noisy aspect previously explained. However, the performance of the model built on CamemBERT-bio, namely CamemBERT-bio + LESA, significantly improves with prefinetuning. This improvement is attributed to the LESA technique, indicating that incorporating Label Embeddings necessitates updating CamemBERT-bio’s embeddings. Essentially, the latent space derived from CamemBERT-bio + LESA evolves to better discriminate between clusters corresponding to each of the eight classes, thereby enhancing the $F_{1}$ score.
TABLE II: Impact of prefinetuning on $F_{1}$ score
Model Without With
Camembert-bio 0.69 0.70
Camembert-bio + LESA 0.79 0.83
VII-2 Impact of Multiple Loss objective
Figure 5: The numbers predicted by Model 2. The $x$ axis represents the ground truth values, and the $y$ axis represents the predicted values. Both axes are displayed on a logarithmic scale
The primary advantage of the number prediction loss function is that it introduces intrinsic magnitude information into the model’s embeddings, an aspect significantly enhanced by incorporating Xval. Through this loss function, we aim for the model to reconstruct numbers with similar magnitudes to the original values. While predicting the exact numbers would lead to memorization, we expect the predictions to be close but not identical. As shown in Figure  5, Model 2 successfully achieves this task.
A detailed examination of the results in Table  III reveals that Model 2 outperforms all other models in nearly every class. Notably, all models struggle with the ventricular Gradient class. This difficulty arises because this class is ambiguous and can be confused with other similar classes, such as aortic gradient, aortic ring, pulmonary gradient, and gradient between earcups. This highlights a limitation of context-based token representation and explains the poor performance of the other models. The range of these gradient values serves as a significant clue for classification, as gradients between earcups are significantly lower than ventricular gradients, for example.
When we subject Model 2 to the same sentence completion test, the results are shown in Figure 6.
Figure 6: Output of a completion task by Model 2. The input sentence is “Patient en détresse respiratoire, gradient VG-VD ad $\langle$ mask $\rangle$ mmgh.”. The figure contains the possible words to fill the mask with their corresponding probabilities. Here the model successfully predict a number which is given by the number prediction head
This improvement was not observed after prefinetuning CamemBERT-bio, as indicated in Table  II. The performance of CamemBERT-bio remains largely unchanged after prefinetuning. Instead, we attribute the improvement to the implementation of multiple objective number representations. By replacing numbers with placeholders, the text—previously unstructured and filled with abbreviations—becomes clearer, making it easier for the model to learn where numbers are expected. Furthermore, incorporating LESA and Xval enhances the model’s ability to more accurately predict the value range of the masked numbers. We aimed to illustrate our claim by comparing the embeddings of a selection of medical terms generated from various models. Since these medical terms are not numbers, we expect to observe the side effects of multiple objective number representations on the rest of the medical vocabulary. We chose cosine similarity as our metric because, as explained in [28], cosine similarity from token embeddings can be random and meaningless unless the model is trained with respect to cosine similarity, which layer normalization may facilitate. Given that this prerequisite is satisfied in our model, we opted for cosine similarity.
We compared the embeddings generated by the pretrained CamemBERT-bio, the finetuned CamemBERT-bio, Model 1, and Model 2. For consistency, we selected the embeddings from the pretrained CamemBERT-bio as the reference. The cosine similarity graphics are displayed in Figure  7. In this figure, each plot represents a cosine similarity matrix $M$ in grayscale. Each row $i$ and column $j$ is indexed by medical terms $\textrm{tok}_{i}$ and $\textrm{tok}_{j}$, respectively. The corresponding cell $M_{i,j}$ contains the cosine similarity score of the embeddings of $\textrm{tok}_{i}$ obtained by the reference model and the embedding of $\textrm{tok}_{j}$ obtained by the model we’re comparing. The darker a cell is, the higher the cosine similarity score.
The second and third matrices in Figure  7 show that the embeddings generated by the finetuned CamemBERT-bio are significantly different from the reference, while those generated by Model 1 (CamemBERT-bio + LESA) are very similar to the reference. This supports the claim from [2] that simple finetuning significantly disrupts the encoded knowledge of pretrained CamemBERT-bio and strongly overfits, while LESA stabilizes the encoded knowledge of pretrained CamemBERT-bio by guiding the self-attention towards tokens with higher affinity. Regarding the impact of multiple objective number representations, the last matrix in Figure  7 shows that the largest embeddings gap with the reference is observed with Model 2. This evidence supports the claim that integrating multiple losses significantly enhances embeddings compared to using only LESA. Unlike the case of the finetuned CamemBERT-bio in the second matrix, the substantial change in embeddings relative to the reference does not indicate overfitting. This assertion is further validated by the higher F1 score obtained on the test dataset.
Figure 7: Evaluation of the impact of multiple objective numbers representation on the words semantic. From left to right we computed the cosine similarity of multiple medical words embeddings generated by pretrained CamemBERT-bio and itself, pretrained CamemBERT-bio and finetuned CamemBERT-bio, pretrained CamemBERT-bio and Model 1, pretrained CamemBERT-bio and Model 2. The darker a cell is, the higher its value
TABLE III: $F_{1}$ score results per class.
O Cp FC D
ELMO 0.99 $\pm$ 0.00 0.76 $\pm$ 0.25 0.62 $\pm$ 0.18 0.68 $\pm$ 0.13
DistilCamemBERT 0.99 $\pm$ 0.00 0.80 $\pm$ 0.16 0.62 $\pm$ 0.09 0.69 $\pm$ 0.14
Camembert-bio 0.99 $\pm$ 0.00 0.78 $\pm$ 0.27 0.68 $\pm$ 0.07 0.52 $\pm$ 0.11
CamemBERT-bio + LESA 0.99 $\pm$ 0.00 0.85 $\pm$ 0.07 0.87 $\pm$ 0.04 0.84 $\pm$ 0.13
CamemBERT-bio + Xval 0.99 $\pm$ 0.00 0.85 $\pm$ 0.14 0.64 $\pm$ 0.07 0.92 $\pm$ 0.06
NumBERT 0.99 $\pm$ 0.00 0.78 $\pm$ 0.10 0.64 $\pm$ 0.04 0.67 $\pm$ 0.15
Model 1 0.99 $\pm$ 0.00 0.85 $\pm$ 0.07 0.87 $\pm$ 0.04 0.84 $\pm$ 0.13
Model 2 0.99 $\pm$ 0.00 0.86 $\pm$ 0.07 0.74 $\pm$ 0.11 0.84 $\pm$ 0.08
SO2 AGPR G CIA/CIV
ELMO 0.83 $\pm$ 0.15 0.87 $\pm$ 0.12 0.24 $\pm$ 0.14 0.81 $\pm$ 0.16
DistilCamemBERT 0.84 $\pm$ 0.05 0.95 $\pm$ 0.02 0.23 $\pm$ 0.02 0.81 $\pm$ 0.15
Camembert-bio 0.85 $\pm$ 0.05 0.93 $\pm$ 0.02 0.22 $\pm$ 0.02 0.80 $\pm$ 0.08
CamemBERT-bio + LESA 0.91 $\pm$ 0.02 0.99 $\pm$ 0.01 0.27 $\pm$ 0.01 0.91 $\pm$ 0.05
CamemBERT-bio + Xval 0.90 $\pm$ 0.03 0.95 $\pm$ 0.01 0.23 $\pm$ 0.15 0.91 $\pm$ 0.08
NumBERT 0.83 $\pm$ 0.07 0.90 $\pm$ 0.13 0.24 $\pm$ 0.10 0.82 $\pm$ 0.11
Model 1 0.91 $\pm$ 0.02 0.99 $\pm$ 0.01 0.27 $\pm$ 0.01 0.91 $\pm$ 0.05
Model 2 0.93 $\pm$ 0.02 0.98 $\pm$ 0.03 0.50 $\pm$ 0.12 0.94 $\pm$ 0.06
VIII Conclusion
This research focused on optimizing numerical understanding by CamemBERT-bio, particularly for classifying numerical values in small medical datasets. We addressed the challenges posed by tokenization and the lack of text structure, which required special handling of numbers. To tackle these issues, we implemented multiple approaches. First, we demonstrated that CamemBERT-bio + LESA could be further improved through prefinetuning with the MLM task, an improvement not seen with CamemBERT-bio alone. Additionally, we developed a multiple-objective numbers representation by combining LESA [2] and Xval [3], resulting in embeddings that incorporate both contextual and magnitude information. Rigorous evaluations of these strategies showed significant enhancements in the performance of CamemBERT-bio.
Acknowledgment
The medical notes utilized in this study were provided by CHU Sainte Justine Hospital and annotated by Dr. Jérôme Rambaud and Dr. Guillaume Sans. The authors express their gratitude to physiotherapist Kevin Albert for his contribution to the literature review on medical benchmarks.
This research was partially supported by the Natural Sciences and Engineering Research Council (NSERC), the Institut de Valorisation des Données de l’Université de Montréal (IVADO), and the Fonds de la Recherche en Santé du Québec (FRQS).